OpenStudy (anonymous):

Solve: (x2 - 3)2 + (x2 - 3) - 2 = 0

OpenStudy (vishweshshrimali5):

\[(x^2-3)^2+x^2-3 - 2 = 0\] Substitute \[x^2-3=y\] Thus eq. becomes \[y^2-y-2 = 0\] \[=> y^2 - 2y+y-2 = 0\] \[=>y(y-2)-1(y-2) = 0\] \[=>(y-1)(y-2)=0\] \[=> y = 1\] or \[y = 2\] \[=> x^2-3x = 1\] or \[x^2- 3x = 2\] Now solve these 2 eq.. to get values pf x

OpenStudy (anonymous):

ok thanks for the help

OpenStudy (vishweshshrimali5):

\[x^2 - 3x -1 = 0\] has 2 roots \[1/2 (3+\sqrt{13})\] \[1/2 (3-\sqrt{13})\] Similarly \[x^2 - 3x - 2 = 0\] has 2 roots \[1/2 (3-\sqrt{17})\] \[1/2 (3+ \sqrt{17})\]

OpenStudy (vishweshshrimali5):

These are your roots